# A magnetic avalanche as the central engine powering a solar flare

### v1

@@
Solar flares are the most powerful, magnetically-driven, explosions in the heliosphere. The nature of magnetic energy release in the solar corona that heats the plasma and accelerates particles in a flare, however, remains poorly understood. Here, we report high-resolution coronal observations of a flare (SOL2024-09-30T23:47) by the Solar Orbiter mission that reveal initially weaker but rapid reconnection events, on timescales of at most a few seconds, leading to a more prominent activity of similar nature that explosively cause a flare. Signatures of this process are further imprinted on the widespread raining plasma blobs with short lifetimes, giving rise to the characteristic ribbon-like emission pattern associated with the flare. Our novel observations unveil the central engine of a flare and emphasize the crucial role of an avalanche-like magnetic energy release mechanism at work.
@@
Flares are powerful explosions produced by the Sun. They can lead to intense electromagnetic radiation, heat plasma to tens of million Kelvin, accelerate particles to relativistic speeds with energies few 10 keV, and propel material into interplanetary space as coronal mass ejections (e.g., Priest & Forbes 2002; Benz 2017). Flares, including the more energetic super-flares, are also common in stars that possess an outer convective zone (Feinstein et al. 2022; Vasilyev et al. 2024).
Almost all the large flares on the Sun (i.e., the M- and X-class events, with peak soft X-ray fluxes W m-2) are caused by a catastrophic destabilization of a filament — a highly-twisted and finely-structured magnetic flux rope with cooler and denser chromospheric gas that is suspended by the magnetic field in the hotter and rarer million Kelvin corona. Such filament eruptions are also thought to be causing stellar flares (Namekata et al. 2021). Processes that destabilize the flux rope and mechanisms for magnetic energy release at flare onset, however, are topics of long-standing debate. Prominent models include magnetohydrodynamic (MHD) processes such as kink and torus instabilities (Hood & Priest 1979; Török et al. 2004; Kliem & Török 2006), magnetic breakout (Antiochos et al. 1999) or a tether-cutting process (Moore et al. 2001).
All these processes must involve reconnection at some stage of the flare evolution, through which magnetic energy is released and converted to other forms (Priest & Forbes 2000). The energy release could also be highly fragmented (Browning et al. 2008; Hood et al. 2009). Previous microwave observations of a large solar flare revealed rapid decay of coronal magnetic field (5 Gauss s-1), that is hypothesized to be caused by turbulent magnetic diffusion (Fleishman et al. 2020). Due to the inherently small spatial scales of a few 100 km and rapid timescales of a few seconds over which impulsive magnetic processes are observed to operate in the corona (Cirtain et al. 2013; Antolin et al. 2021; Berghmans et al. 2021; Chitta et al. 2022), high (spatial and temporal) resolution coronal observations are crucial to capture magnetic dynamics, as traced by plasma structures, and to investigate various aspects of the flare process.
@@ 2 High-resolution observations of a solar flare
Here we report high-spatial-resolution (210 km) and very-high-cadence (2 s) extreme ultraviolet (EUV) observations of both the preflare and impulsive phases of a major M7.7-class flare on 2024 September 30, which started around universal time as measured at Earth (UT) 23:47, SOL2024-09-30T23:47, (Fig. 5), caught by the Extreme Ultraviolet Imager (EUI; Rochus et al. 2020) onboard the Solar Orbiter spacecraft (Müller et al. 2020). These are among the highest resolution EUV observations of a flare ever recorded, thanks to the EUV high-resolution imager (HRIEUV) of the EUI telescope (Figs. 1 and 6; movies S1 and S2). The instrument samples emission in the 174 Å range from 1 MK coronal gas due to contributions from the Fe ix (at 171.1 Å) and Fe x (at 174.5 Å and 177.2 Å) spectral lines. The event was also captured by the Spectrometer/Telescope for Imaging X-rays (STIX; Krucker et al. 2020), the Spectral Imaging of the Coronal Environment (SPICE; SPICE Consortium et al. 2020) instrument, and the Polarimetric and Helioseismic Imager (SO/PHI; Solanki et al. 2020), three other remote-sensing instruments on Solar Orbiter that sample different layers and different temperature regimes of the solar atmosphere. Thus, our novel observations offer a comprehensive testbed to understand the flare onset and its impact on different layers of the solar atmosphere (complete observational details, including information on the flare location on the Sun, and the type of magnetic environment in which it occurred, are provided in Appendix A).
@@ 3 Reconnection throughout the flaring process
To give a brief overview of the flare, it was associated with an unstable flux rope containing a filament that initially appeared in our HRIEUV images as a darker feature suspended in the corona due to EUV absorption by the embedded chromospheric material (Fig. 6). The filament can also be seen at temperatures below 1 MK, sampled by SPICE spectroscopic observations (Fig. 7 and the accompanying online animation). STIX recorded an impulsive enhancement of the hard X-ray (HXR) emission due to bremsstrahlung with energies in excess of 25 keV that lasted for about 2 minutes between UT 23:46 and UT 23:48 (Figs. 8 and 9). The following description of events is with respect to this epoch of the impulsive HXR phase around 23:47 UT that marks the flare onset, when both the soft and hard X-ray emission rise steeply.
@@
A distinct X-shaped configuration develops in the atmosphere, connecting the dark filament threads to the adjacent complex loop system (Fig. 6) during the preflare phase, some 30 minutes prior to the impulsive phase. Such an X-shaped structure is a typical observational signature of the various 3D geometries in which magnetic reconnection can occur (Pontin & Priest 2022). Here we observed persistent formation and bi-directional propagation of new bright strands at the site of the X-shaped structure. We demonstrate the spatial and temporal complexity of this evolution over two short 32 s-long intervals in Fig. 2, using sequences of running-difference images.
@@
There are two distinct properties of this evolution that stand out. Firstly, these dynamics are so rapid that we see new strands developing in almost every snapshot of our observations. Secondly, the formation of strands closely resembles that of winding or unwinding of the loops (Fig. 2A, Movie S3). As this X-shaped configuration continues to evolve, intensity variations and enhancements become more rapid as well, with changes being recorded at the level of the observational cadence (Fig. 2B, Movie S4). At the same time, the newly formed strands are bi-directionally swept away with plane-of-sky speeds of 100–200 km s-1, slower in the initial stages, but tending to be faster closer to the impulsive phase of the flare. Based on these properties, we interpret these dynamics to be associated with magnetic reconnection near the X-shaped structure.
@@
Despite this reconfiguration of the magnetic field and the concurrent EUV enhancement during this period, up until UT 23:40, we did not detect any clear enhancement of HXR signal in the STIX data, indicating the absence of both higher temperature plasma 10 MK and accelerated non-thermal particles that are required to produce the HXR emission in the first place. In comparison to the overall further development of signatures in the UV to X-ray wavelengths, the reconnection dynamics in this early pre-impulsive phase of the flare are deemed weaker in nature.
@@
From the perspective of coronal dynamics as captured by HRIEUV, however, there is a continued increase in the spatial complexity surrounding the X-shaped configuration. This is further reflected in the evolution of the adjacent filament. In the initial stages (time period around sequence-A in Fig. 2A), despite some minor morphological changes, the filament remains nearly unperturbed (Movie S2). But the more active stage covered by sequence-B initiates the rise of the filament as the latter continues to separate from the complex loop system. This separation creates a flurry of rapidly changing EUV signatures along with complex bi-directional flow patterns at the interface of the two features. The lower resolution SPICE spectral maps do not capture this complex evolution. Nevertheless, they reveal the presence of multi-thermal plasma in the observed range of 104–106 K (Movie S6). Furthermore, we detected the increase in the STIX HXR signal to be co-spatial with this interface region around UT 23:44 (Movie S1), indicating higher plasma temperatures than before, further implying continued increase of energy transfer from the magnetic field to plasma through reconnection.
@@
We observed that the rise of the filament is marked by prominent reconnection along its length, primarily at three specific locations. In Fig. 3 we display this evolution along these three locations (panels A–C) that are characterized by increased speeds in the reconnection flows. Bi-directional flows with high speeds over 400 km s-1 are observed at the site of complete disconnection of the filament from the adjacent loop system. After this disconnection, the filament, still anchored to the Sun at the other end, exhibits rapid unwinding motions, with projected speeds over 250 km s-1 (panels D–E). The EUV emission at the filament footpoint begins to increase at the same time as the reconnection event in panel E, both from the perspective of HRIEUV and the Atmospheric Imaging Assembly (AIA; Lemen et al. 2012) on board the near-Earth Solar Dynamics Observatory (SDO; Pesnell et al. 2012, more details provided in Appendix A). Higher energy HXR emission in the 25–50 keV, which is consistent with bremsstrahlung caused by accelerated particles depositing energy in the higher dense footpoint region (Fig. 9) begins to increase nearly a minute later, lagging the intense EUV emission at the site of reconnection by at least 10 s.
@@
Overall, these filament dynamics are more discrete yet more energetic than the evolution that we already observed in the adjacent loop system. Nevertheless, both share similarities in that they clearly evolve on short timescales, limited only by our observational cadence of 2 s. STIX observations reveal that while the location of the filament disconnection and the complex loop system is co-spatial to lower energy HXR sources below 14 keV, the side of the filament that is still rooted to the Sun exhibited both low and high energy HXR sources more than a minute later (compare panels C and E in Fig. 3).
@@ 4 Implications for the flare onset – a magnetic avalanche
A combination of emergence of magnetic flux through the atmosphere, its convergence and cancellation near the polarity inversion line, coupled with surface shear motions, are thought to load magnetic energy into a magnetic flux rope containing a filament, which is eventually released as a flare through an instability (van Ballegooijen & Martens 1989; Chen & Shibata 2000; Chintzoglou et al. 2019; Török et al. 2024). These processes are closely governed by convective motions that evolve on timescales of minutes at a granular scale (Hirzberger et al. 1999). Although we cannot determine the distribution of magnetic energy release through time (due to the lack of coronal magnetic field observations in our case), our novel HRIEUV observations suggest that reconnection is progressing on timescales as short as 2 s.
@@
Because of such a large timescale difference between the surface convection and atmospheric reconnection, photospheric driver and the coronal flaring region will be dynamically decoupled. This decoupling can be seen as continual breakage of loop emission patterns at the site of the X-shaped configuration (Movie S2). Despite this, there appears to be a clear ordering from initially weak reconnection events to progressively more energetic ones according to our EUV and HXR observations. This progression is observed to manifest also in the flux rope containing the filament itself, in which the reconnection signatures are initiated between a few threads around UT 23:46, which rapidly propagate to the whole body of the filament in a matter of 2 minutes. Based on this, we suggest that the magnetic energy release evolves through an avalanche-like self-organized criticality process.
@@
Simple avalanche concepts have been invoked previously to explain the statistical power-law distribution of HXR peak fluxes from a large number of flares (Bak et al. 1987; Lu & Hamilton 1991; Charbonneau et al. 2001; Aschwanden et al. 2016). These concepts are further extended to explain the power-law frequency distribution of flares from a large sample of main sequence stars (Feinstein et al. 2022). But these statistics do not capture the physics of underlying individual flares. A full MHD model of avalanche behavior shows how the instability and reconnection of one magnetic thread or strand (i.e., a tiny elementary flux tube) could cause neighboring threads to go unstable in a MHD avalanche (Hood et al. 2016; Reid et al. 2023). Our observations are consistent with this scenario. They reveal how reconnection avalanches occurring in different parts of the system (the complex loop system and the main erupting flux rope itself) at different stages evolve cohesively to produce the observed characteristics of the flare. This mechanism, captured by our observations, offers a natural explanation to the universality of flare energy power-law distribution in the main sequence stars.
@@ 5 Implications for particle acceleration and energy transport
Reconnection timescales of about 2 s that we observed here closely match flare magnetic reconnection models designed to explain HXR emission (Fletcher & Martens 1998). In the standard solar flare model (Carmichael 1964; Sturrock 1966; Hirayama 1974; Kopp & Pneuman 1976), including magnetic breakout (Antiochos et al. 1999) and tether-cutting runaway reconnection models (Moore et al. 2001), a current sheet forms below the center of the slowly rising filament, where reconnection and particle acceleration are thought to ensue. In this scenario, reconnection outflows with super-magnetosonic speeds emerging from the current sheet may also produce a termination shock that accelerates particles at the apexes of flaring loops (Chen et al. 2015). In our event, such a process occurs weakly during the preflare phase.
@@
As shown in Fig. 1, high energy HXR emission originates from a compact region at one footpoint of the erupting flux rope, which suggests that the reconnection is occurring within the flux rope rather than below it. At the same time, the lower energy components have elongated structures aligned with the leg of the filament itself (additional discussion on various HXR sources is presented in Appendix B). This location is directly below the location where we observed unwinding filament threads with rapid flows (Fig. 3, 6). At this time, almost all the filament material is completely disconnected from the adjacent complex loop system where we observed weaker reconnection events initially. The filament itself is rooted in a unipolar region adjacent to a sunspot (Figs. 8 and 10) and so we do not expect the current sheet as in the standard solar flare model to be dominating energy release in the impulsive phase. Instead, the observations suggest that reconnection within the erupting flux rope is able to accelerate non-thermal particles to energies beyond 18 keV (Fig. 9), producing the HXR emission. Indeed, HXR emission at the feet of erupting filaments has been previously seen (Liu et al. 2009; Stiefel et al. 2023). But due to the lack of suitable observations, the evolution and internal dynamics of the erupting filament could not be evaluated in those studies. Our observations thus demonstrate that the commonly invoked scenario of particle acceleration in solar flares is not the whole picture, and that magnetic avalanches within an erupting magnetic flux rope should be considered instead for the impulsive phase.
@@
A concurrent brightening of the dense filament with the unwinding motions (Fig. 3) suggests that the released magnetic energy is locally deposited to heat the plasma. Such unwinding processes have been observed to produce collimated jets and localized brightenings in some filamentary structures and coronal loops (Cirtain et al. 2013; Antolin et al. 2021; Chitta et al. 2022). It is likely that the filament plasma is heated directly rather than being supplied by chromospheric ablation (i.e., the process that explains the heating of loops in the main phase of the flare). This is because such an ablation would have to supply hot material to fill the filament up to the projected heights of 20 Mm in about 20 s, implying speeds of more than a 1000 km s-1 in this case. But such high speeds of upflows into the loops have not been observed. Such an ablation process would play a role later in the main phase of the flare as the flare arcade develops a few minutes after the impulsive phase (Movie S2).
@@
Once the magnetic energy is released in the corona, it will be transported along the field lines and be deposited in the lower atmosphere. In our HRIEUV observations, we identified a stream of raining plasma blobs at the footpoints of the reconnecting loop system (Fig. 4). An elongated thread – bright in the EUV emission – appears first, and is then modulated by a train of small-scale plasma blobs with spatial sizes at the resolution limit of our observations and that live for at most a few seconds. The thread itself lives for about 10–20 s. However, there is no spatial or temporal coherence of the raining plasma blobs among neighboring threads despite their separation being only a few 100 km. These threads in combination appear as flare ribbons. Within the flaring region itself (Fig. 6), these raining plasma blobs are initiated at least 5 minutes prior to the strong HXR impulse and continue to persist several minutes after the impulsive phase has ceased (Movie S5). The same behavior is also seen at remote locations from at least 4 minutes prior to the impulsive phase, some 100 Mm away from the flare site, that are connected via long coronal loops (Figs. 6, 10, 12, 13, and Movie S2). To our knowledge, this is the first time we have been able to identify such a rapidly evolving nature of flare ribbons in the EUV. Overall, the timescales of these raining plasma blobs compare well with the avalanche-type energy release that we inferred from the EUV signatures of coronal reconnection. We cannot determine the nature of these plasma blobs with these observations, but their commonality among different near and remote locations from the flare site over a long duration provides constraints on the dominant energy transport process at work in flares.

### v2

@@
Solar flares are the most powerful, magnetically driven, explosions in the heliosphere. The nature of magnetic energy release in the solar corona that heats the plasma and accelerates particles in a flare, however, remains poorly understood. Here, we report high-resolution coronal observations of a flare by the Solar Orbiter mission that reveal initially weaker but rapid reconnection events, on timescales of a few seconds at most, leading to a more prominent activity of a similar nature that explosively causes a flare. Signatures of this process are further imprinted on the widespread raining plasma blobs with short lifetimes, giving rise to the characteristic ribbon-like emission pattern associated with the flare. Our observations unveil the central engine of a flare and emphasize the crucial role of an avalanche-like magnetic energy release mechanism at work.
@@
Flares are powerful explosions produced by the Sun. They can lead to intense electromagnetic radiation, heat plasma to tens of million Kelvin, accelerate particles to relativistic speeds with energies greater than a few 10 keV, and propel material into interplanetary space as coronal mass ejections (e.g., Priest & Forbes 2002; Benz 2017). Flares, including the more energetic super-flares, are also common in stars that possess an outer convective zone (Feinstein et al. 2022; Vasilyev et al. 2024).
Almost all the large flares on the Sun (i.e., the M- and X-class events, with peak soft X-ray fluxes W m-2) are caused by a catastrophic destabilization of a filament — a highly twisted and finely structured magnetic flux rope with cooler and denser chromospheric gas that is suspended by the magnetic field in the hotter and rarer million Kelvin corona. Such filament eruptions are also thought to cause stellar flares (Namekata et al. 2021). Processes that destabilize the flux rope and mechanisms for magnetic energy release at flare onset, however, are topics of long-standing debate. Prominent models include magnetohydrodynamic (MHD) processes such as kink and torus instabilities (Hood & Priest 1979; Török et al. 2004; Kliem & Török 2006), magnetic breakout (Antiochos et al. 1999), or a tether-cutting process (Moore et al. 2001).
All of these processes must involve reconnection at some stage of the flare evolution, through which magnetic energy is released and converted to other forms (Priest & Forbes 2000). The energy release could also be highly fragmented (Browning et al. 2008; Hood et al. 2009). Previous microwave observations of a large solar flare revealed rapid decay of the coronal magnetic field (5 Gauss s-1), which is hypothesized to be caused by turbulent magnetic diffusion (Fleishman et al. 2020). Due to the inherently small spatial scales of a few 100 km and rapid timescales of a few seconds over which impulsive magnetic processes are observed to operate in the corona (Cirtain et al. 2013; Antolin et al. 2021; Berghmans et al. 2021; Chitta et al. 2022), high (spatial and temporal) resolution coronal observations are crucial to capture magnetic dynamics, as traced by plasma structures, and to investigate various aspects of the flare process.
@@ 2 High-resolution observations of a solar flare
Here we report high-spatial-resolution (210 km) and very-high-cadence (2 s) extreme-ultraviolet (EUV) observations of both the pre-flare and impulsive phases of a major M7.7-class flare on 2024 September 30, which started around universal time as measured at Earth (UT) 23:47, SOL2024-09-30T23:47 (Fig. 6), and was captured by the Extreme Ultraviolet Imager (EUI; Rochus et al. 2020) on board the Solar Orbiter spacecraft (Müller et al. 2020). These are among the highest-resolution EUV observations of a flare ever recorded, thanks to the EUV high-resolution imager (HRIEUV) of the EUI telescope (Figs. 1 and 2; movies S1 and S2). The instrument samples emission in the 174 Å range from 1 MK coronal gas due to contributions from the Fe ix (at 171.1 Å) and Fe x (at 174.5 Å and 177.2 Å) spectral lines. The event was also captured by the Spectrometer/Telescope for Imaging X-rays (STIX; Krucker et al. 2020), the Spectral Imaging of the Coronal Environment (SPICE; SPICE Consortium et al. 2020) instrument, and the Polarimetric and Helioseismic Imager (SO/PHI; Solanki et al. 2020), three other remote-sensing instruments on Solar Orbiter that sample different layers and different temperature regimes of the solar atmosphere. Thus, our novel observations offer a comprehensive test bed to understand the flare onset and its impact on different layers of the solar atmosphere. Complete observational details, including information on the flare location on the Sun, and the type of magnetic environment in which it occurred, are provided in Appendix A.
@@ 3 Reconnection throughout the flaring process
To give a brief overview of the flare, it was triggered by a complex loop system exhibiting an X-shaped morphology (Fig. 2) that destabilized a flux rope containing a filament. This filament initially appeared in our HRIEUV images as a darker feature suspended in the corona due to EUV absorption by the embedded chromospheric material (Fig. 2). The filament can also be seen at temperatures below 1 MK, sampled by SPICE spectroscopic observations (Fig. 7 and the accompanying online animation). STIX recorded an impulsive enhancement of the hard X-ray (HXR) emission due to bremsstrahlung with energies in excess of 25 keV that lasted for about 2 minutes between UT 23:46 and UT 23:48 (Figs. 8 and 9). The following description of events is with respect to this epoch of the impulsive HXR phase around UT 23:47 that marks the flare onset, when both the soft and HXR emission rise steeply.
@@
A distinct X-shaped configuration develops in the atmosphere, connecting the dark filament threads to the adjacent complex loop system (movie associated with Fig. 2) during the pre-flare phase, some 30 minutes prior to the impulsive phase. Such an X-shaped structure is a typical observational signature of the various 3D geometries in which magnetic reconnection can occur (Pontin & Priest 2022). Here we observed persistent formation and bi-directional propagation of new bright strands at the site of the X-shaped structure. We demonstrate the spatial and temporal complexity of this evolution over two short 32 s long intervals in Fig. 3, using sequences of running-difference images.
@@
There are two distinct properties of this evolution that stand out. Firstly, these dynamics are so rapid that we see new strands developing in almost every snapshot of our observations. Secondly, the formation of strands closely resembles that of the winding or unwinding of the loops (Fig. 3A, Movie S3). As this X-shaped configuration continues to evolve, intensity variations and enhancements become more rapid as well, with changes being recorded at the level of the observational cadence (Fig. 3B, Movie S4). At the same time, the newly formed strands are bi-directionally swept away with plane-of-sky speeds of 100–200 km s-1, which are slower in the initial stages but they tend to be faster closer to the impulsive phase of the flare. It could be argued that the fine-scale dynamics that the EUI observed in this pre-flare phase are simply a reflection of any nonuniformity of the magnetic structure coupled with temperature and density fluctuations. A strong localized fluctuating heat source is, however, required to generate these intense thermodynamic variations. Here, the rapid development of magnetic strands on timescales down to 2 s, combined with the detection of their bi-directional separation and localized strong intensity enhancements, all point to the small-scale reconnection scenario as the underlying process fueling the dynamics around the X-shaped configuration.
@@
Despite this reconfiguration of the magnetic field and the concurrent EUV enhancement during this period, up until UT 23:40, the high-energy HXR signal in the STIX data is faint. This indicates a weak presence, if any, of accelerated nonthermal particles. In comparison to the overall further development of signatures in the UV to X-ray wavelengths, the reconnection dynamics in this early pre-impulsive phase of the flare are deemed weaker in nature.
@@
From the perspective of coronal dynamics as captured by HRIEUV, however, there is a continued increase in the spatial complexity surrounding the X-shaped configuration. This is further reflected in the evolution of the adjacent filament. In the initial stages (the time period around sequence-A in Fig. 3A), despite some minor morphological changes, the filament remains nearly unperturbed (Movie S2). However, the more active stage covered by sequence B initiates the rise of the filament as the latter continues to separate from the complex loop system. This separation creates a flurry of rapidly changing EUV signatures along with complex bi-directional outflow patterns at the interface of the two features. Given the broad thermal response of the HRIEUV instrument, it is uncertain what plasma temperature it is probing. Here we make use of SPICE observations to mitigate this uncertainty. Though the lower-resolution SPICE spectral maps do not capture this complex evolution, they nevertheless reveal the presence of multi-thermal plasma in the observed range of 104–106 K (Movie S6). Furthermore, we detected the increase in the STIX HXR signal to be co-spatial with this interface region around UT 23:44 (Movie S1), indicating higher plasma temperatures than before. This further implies a continued increase of energy transfer from the magnetic field to plasma through reconnection.
@@
We observed that the rise of the filament is marked by prominent reconnection along its length, primarily at three specific locations. In Fig. 4 we display this evolution along these three locations (panels A–C) that are characterized by increased speeds in the reconnection outflows. Bi-directional flows with high speeds over 400 km s-1 are observed at the site of complete disconnection of the filament from the adjacent loop system. After this disconnection, the filament, which is still anchored to the Sun at the other end, exhibits rapid unwinding motions, with projected speeds over 250 km s-1 (panels D–E). The EUV emission at the filament footpoint begins to increase at the same time as the reconnection event in panel E, both from the perspective of HRIEUV and the Atmospheric Imaging Assembly (AIA; Lemen et al. 2012) on board the near-Earth Solar Dynamics Observatory (SDO; Pesnell et al. 2012, more details are provided in Appendix A). Higher-energy HXR emission in the 25–50 keV, which is consistent with bremsstrahlung caused by accelerated particles depositing energy in the higher, dense footpoint region (Fig. 9), begins to increase nearly a minute later, lagging the intense EUV emission at the site of reconnection by at least 10 s.
@@
Overall, these filament dynamics are more discrete, yet more energetic, than the evolution that we already observed in the adjacent loop system. Nevertheless, both share similarities in that they clearly evolve on short timescales of 2 s that we could resolve in our observations. STIX observations reveal that while the location of the filament disconnection and the complex loop system is co-spatial to lower-energy HXR sources below 14 keV, the side of the filament that is still rooted to the Sun exhibited both low- and high-energy HXR sources more than a minute later (compare panels C and E in Fig. 4).
@@ 4 Implications for the flare onset – A magnetic avalanche
A combination of the emergence of magnetic flux through the atmosphere, its convergence, and cancellation near the polarity inversion line, coupled with surface shear motions, are thought to load magnetic energy into a magnetic flux rope containing a filament, which is eventually released as a flare through an instability (van Ballegooijen & Martens 1989; Chen & Shibata 2000; Chintzoglou et al. 2019; Török et al. 2024). These processes are closely governed by convective motions that evolve on timescales of minutes on a granular scale (Hirzberger et al. 1999). Although we cannot determine the distribution of magnetic energy release through time (due to the lack of coronal magnetic field observations in our case), our novel HRIEUV observations suggest that reconnection is progressing on timescales as short as 2 s.
@@
Because of such a large timescale difference between the surface convection and atmospheric reconnection, the photospheric driver and coronal flaring region will be dynamically decoupled. This decoupling can be seen as continual breakage of loop emission patterns at the site of the X-shaped configuration (Movie S2). Despite this, there appears to be a clear ordering from initially weak reconnection events to progressively more energetic ones according to our EUV and HXR observations. This progression is observed to manifest also in the flux rope containing the filament itself, in which the reconnection signatures are initiated between a few threads around UT 23:46, which rapidly propagate to the whole body of the filament in a matter of 2 minutes. Based on this, we suggest that the magnetic energy release evolves through an avalanche-like self-organized criticality process. As our observations reveal, this process is able to generate plasma heating over a wide range of temperatures (as captured by the SPICE, EUI, and STIX instruments).
@@
Simple avalanche concepts have been invoked previously to explain the statistical power-law distribution of HXR peak fluxes from a large number of flares (Bak et al. 1987; Lu & Hamilton 1991; Charbonneau et al. 2001; Aschwanden et al. 2016). These concepts are further extended to explain the power-law frequency distribution of flares from a large sample of main sequence stars (Feinstein et al. 2022). However, these statistics do not capture the physics underlying individual flares. A full MHD model of avalanche behavior shows how the instability and reconnection of one magnetic thread or strand (i.e., a tiny elementary flux tube) could cause neighboring threads to become unstable in a MHD avalanche (Hood et al. 2016; Reid et al. 2023). Additional discussion on the universality of magnetic avalanche reconnection and the current state of MHD simulations is presented in Appendix C. Our observations are consistent with this scenario. They reveal how reconnection avalanches occurring in different parts of the system (the complex loop system and the main erupting flux rope itself) at different stages evolve cohesively to produce the observed characteristics of the flare. This mechanism, which was captured by our observations, offers a natural explanation to the universality of the flare energy power-law distribution in the main sequence stars. We suggest that such a rapid small-scale reconnection in and around the flaring flux rope can also mimic the effects of turbulent magnetic diffusion (Fleishman et al. 2020).
@@ 5 Implications for particle acceleration and energy transport
Reconnection timescales of about 2 s that we observed here closely match flare magnetic reconnection models designed to explain HXR emission (Fletcher & Martens 1998). In the standard solar flare model (Carmichael 1964; Sturrock 1966; Hirayama 1974; Kopp & Pneuman 1976), including magnetic breakout (Antiochos et al. 1999) and tether-cutting runaway reconnection models (Moore et al. 2001), a current sheet forms below the center of the slowly rising filament, where reconnection and particle acceleration are thought to ensue. In this scenario, reconnection outflows with super-magnetosonic speeds emerging from the current sheet may also produce a termination shock that accelerates particles at the apexes of flaring loops (Chen et al. 2015). In relation to this standard solar flare picture, however, our observations show only weak reconnection events underneath the rising filament during the pre-flare phase. Nevertheless, the pre-flare phase activity does support the tether-cutting magnetic reconnection model (Moore et al. 2001).
@@
As shown in Fig. 1, high-energy HXR emission originates from a compact region at one footpoint of the erupting flux rope, which suggests that the reconnection is occurring within the flux rope rather than below it. At the same time, the lower-energy components have elongated structures aligned with the leg of the filament itself (an additional discussion on various HXR sources is presented in Appendix B). This location is directly below the location where we observed unwinding filament threads with rapid flows (Fig. 4, 2). At this time, almost all the filament material is completely disconnected from the adjacent complex loop system where we observed weaker reconnection events initially. The filament itself is rooted in a unipolar region adjacent to a sunspot (Figs. 8 and 10) and thus we do not expect the current sheet as in the standard solar flare model to dominate energy release in the impulsive phase. Instead, the observations suggest that reconnection within the erupting flux rope is able to accelerate nonthermal particles to energies beyond 18 keV (Fig. 9), producing the HXR emission along with photospheric imprints. Indeed, HXR emission at the feet of erupting filaments has been previously seen (Liu et al. 2009; Stiefel et al. 2023). However, due to the lack of suitable observations, the evolution and internal dynamics of the erupting filament could not be evaluated in those studies. Our observations thus demonstrate that the commonly invoked scenario of particle acceleration within a current sheet below the erupting filament as in a standard solar flare is incomplete. Specifically, particle acceleration caused by a magnetic avalanche reconnection within an erupting magnetic flux rope could also play a role in the impulsive phase.
@@
A concurrent brightening of the dense filament with the unwinding motions (Fig. 4) suggests that the released magnetic energy is locally deposited to heat the plasma. Such unwinding processes have been observed to produce collimated jets and localized brightenings in some filamentary structures and coronal loops (Cirtain et al. 2013; Antolin et al. 2021; Chitta et al. 2022). It is likely that the filament plasma is heated directly rather than being supplied by chromospheric ablation (i.e., the process that explains the heating of loops in the main phase of the flare). This is because such an ablation would have to supply hot material to fill the filament up to the projected heights of 20 Mm in about 20 s, implying speeds of more than a 1000 km s-1 in this case. However, such high speeds of upflows into the loops have not been observed. Such an ablation process would play a role later in the main phase of the flare as the flare arcade develops a few minutes after the impulsive phase (Movie S2).
@@
Once the magnetic energy is released in the corona, it will be transported along the field lines and be deposited in the lower atmosphere. In our HRIEUV observations, we identified a stream of raining plasma blobs at the footpoints of the reconnecting loop system, forming a sequence of threads (Fig. 5). When analyzed individually, an elongated thread, which is bright in the EUV emission, appears first. It is then modulated by a train of small-scale plasma blobs with spatial sizes at the resolution limit of our observations and that live for a few seconds at most. Each thread itself lives for about 10–20 s. However, there is no spatial or temporal coherence of the raining plasma blobs among neighboring threads despite their separation being only a few 100 km. Collectively, these threads appear as flare ribbons. Within the flaring region itself (Fig. 2), these raining plasma blobs are initiated at least 5 minutes prior to the strong HXR impulse and continue to persist several minutes after the impulsive phase has ceased (Movie S5). The same behavior is also seen at remote locations at least 4 minutes prior to the impulsive phase, some 100 Mm away from the flare site; they are connected via long coronal loops (Figs. 2, 10, 12, 13, and Movie S2). The identification of such a rapidly evolving nature of flare ribbons in the EUV is possible due to the high-resolution HRIEUV data. The timescales of these raining plasma blobs compare well with the avalanche-type energy release that we inferred from the EUV signatures of coronal reconnection.
Chromospheric and coronal blob-like features ejecting from sites of magnetic reconnection are commonly observed in high-resolution observations (e.g., Yan et al. 2022; Cheng et al. 2023). Given their spatial proximity to the reconnecting plasma sheets, such blobs are interpreted as plasmoids that form in the current sheets due to the secondary tearing instability (Yan et al. 2022). In our case, the observed blobs occur along the ribbons in the transition region that are far from the presumed reconnection site in the corona. Should the observed blobs be plasmoids, then it is unlikely that they retain their compact identity as they are transported along the newly reconnection magnetic loops. Therefore, we suggest that the blobs are a result of the impulsive injection of particles from the reconnection sites and not the plasmoids themselves. These high-energy particles could thermalize in the denser transition region at the base of reconnecting loops, rapidly heating the plasma (Testa et al. 2014) and resulting in the localized blob-like emission enhancements that we observed. The existence of raining plasma blobs along with the spatial and temporal disconnection neighboring threads, as seen among different near and remote ribbon locations from the flare site over a long duration, provides constraints on the dominant energy transport process at work in flares.